### v1

# HySem: A context length optimized LLM pipeline for unstructured tabular extraction

Abstract
Regulatory compliance reporting in the pharmaceutical industry relies on detailed tables, but these are often under-utilized beyond compliance due to their unstructured format and arbitrary content. Extracting and semantically representing tabular data is challenging due to diverse table presentations. Large Language Models (LLMs) demonstrate substantial potential for semantic representation, yet they encounter challenges related to accuracy and context size limitations, which are crucial considerations for the industry applications. We introduce HySem, a pipeline that employs a novel context length optimization technique to generate accurate semantic JSON representations from HTML tables. This approach utilizes a custom fine-tuned model specifically designed for cost- and privacy-sensitive small and medium pharmaceutical enterprises. Running on commodity hardware and leveraging open-source models, our auto-correcting agents rectify both syntax and semantic errors in LLM-generated content. HySem surpasses its peer open-source models in accuracy and provides competitive performance when benchmarked against OpenAI GPT-4o and effectively addresses context length limitations, which is a crucial factor for supporting larger tables.
HySem: A context length optimized LLM pipeline for unstructured tabular extraction
Narayanan PP1 Indian Institute of Science, Bangalore pnkrishnan27@gmail.com                        Anantharaman Palacode Narayana Iyer1 JNResearch Labs LLP, Bangalore ananth@jnresearch.com
1 Introduction
Tables are a fundamental method for presenting structured information across various industrial applications, including regulatory information in the pharmaceutical industry and financial reports for businesses. Some shared characteristics and inherent challenges in tables as mentioned by Fang et al. (2024) include Heterogeneity Borisov et al. (2023), Sparsity, Context-based interconnection Liu et al. (2023), Lack of prior knowledge Borisov et al. (2023), Borisov et al. (2024), etc. Recent advancements in Large Language Models (LLMs) have shown promise in addressing these challenges for various table related tasks. Some table related tasks and techniques include entity matching and data-imputation Li et al. (2023), Korini and Bizer (2023), tabular Q&A Chen et al. (2024), schema augmentation Narayan et al. (2022), Serialization Yin et al. (2020) Yan et al. (2024), Table Manipulation Zhang et al. (2024), Table Understanding Lobo et al. (2023), Prompt Engineering Zheng et al. (2023), Table RAG Sundar and Heck (2023). Role-play Zhao et al. (2023).
In typical companies, the vast knowledge contained within these tables often remains untapped for actionable insights due to the lack of effective tools and methodologies for extracting and interpreting the tabular data. For instance, documents related to regulatory compliance, such as Annual Product Quality Reviews (APQR), fulfill compliance requirements but are rarely utilized beyond their initial purpose. To leverage the abundant resources of knowledge present in tabular form and stored in various formats like PDF, Word or HTML it is essential to have an accurate extraction system that transforms the content in to a semantically meaningful format. Several works have investigated the LLM’s performance for different input tabular formats such as JSON formats Singha et al. (2023), HTML tables Sui et al. (2024), Markdown Narayan et al. (2022). Accurate semantic representations of tables in a structured form allows automated processing and that can significantly aid the downstream processes, particularly those that use Large Language Models (LLM) as a key component of their workflows. Extracting complex tabular data from semi-structured sources and providing a semantic representation is a hard problem due to the lack of any standards in tabular data representations. This results in a huge diversity and arbitrariness in the way real world tables are presented. The language understanding ability of LLMs and their proven results present an immense opportunity to develop semantics driven represen-
Figure 1: HySem Pipeline Architecture.
tations. However, LLM are prone to hallucinations Huang et al. (2023) Akhtar et al. (2023), the generated outputs may have accuracy challenges and the output formatting may even have syntax errors. In a number of cases, even state of the art LLMs are not very accurate in capturing the complete tabular information.
In this paper, we propose HySem, a context-optimized pipeline designed to generate accurate semantic JSON representations from HTML tables using a custom fine-tuned model. Our research goals were formulated in accordance with industry needs, particularly for small and medium pharmaceutical enterprises, where cost and privacy considerations are paramount.
1.1 Key Goals
  1. 1.
    On-Premise Model Deployment: Models should be run on-premise due to data security considerations.
  2. 2.
    Commodity Hardware Compatibility: Models should operate on commodity hardware with widely available GPUs.
  3. 3.
    Model Size Limitation: The model size should be less than 10 billion parameters to ensure compatibility with GPUs supporting 16 GB RAM.
  4. 4.
    Open-Source Software Stack: The software stack should leverage open-source tools and models.
  5. 5.
    Maximized Extraction Accuracy: Extraction accuracy should be maximized, as any inaccuracies would necessitate costly human correction efforts.
Guided by the above goals, our contributions include:
  • Context Optimizer: A novel context optimizer that rewrites input text to significantly reduce the token count.
  • Semantic Synthesizer: A custom fine-tuned model that generates accurate semantic representations of HTML tables.
  • Automated Syntax Correction: An automated correction system to address syntax errors.
  • Evaluation Methodology: A novel evaluation methodology that measures the performance of our pipeline using both intrinsic and extrinsic metrics.
Input tokens Tokenizer Number of tokens
IRBESARTAN 75MG FILM COATED TABLETS ’IR’, ’B’, ’ES’, ’ART’, ’AN’, ’ ’, ’75’, ’MG’, ’ FIL’, ’M’, ’ CO’, ’ATED’, ’ TABLE’, ’TS’ 14
PANTOPRAZOLE 20 MG GASTRO - RESISTANT TABLETS ’P’, ’ANT’, ’OP’, ’RA’, ’Z’, ’OLE’, ’ ’, ’20’, ’ MG’, ’ G’, ’AST’, ’RO’, ’ -’, ’ RES’, ’IST’, ’ANT’, ’ TABLE’, ’TS’ 18
MPR:PNAHG:A:UK01/R6 ’MP’, ’R’, ’:P’, ’NA’, ’HG’, ’:A’, ’:’, ’UK’, ’01’, ’/R’, ’6’ 11
ML06:FPS:I:INAEG:AU01-002 ’ML’, ’06’, ’:’, ’FPS’, ’:I’, ’:’, ’INA’, ’EG’, ’:’, ’AU’, ’01’, ’-’, ’002’ 13
0619R11486 ’061’, ’9’, ’R’, ’114’, ’86’ 6
08/02/20 ’08’, ’/’, ’02’, ’/’, ’20’ 5
Table 1: Key terms used in a typical compliance document filed by the pharmaceutical industry and the corresponding tokens.
2 Method
Our proposed LLM Pipeline, HySem, aims to address the challenge of converting unprocessed raw HTML tables into a semantic JSON data structure while addressing the limitations of LLMs concerning context length and inference time. By adopting a novel strategy for optimizing the number of tokens, we aim to meet industry standards for high accuracy and reduced inference time, particularly for regulatory compliance reporting. Our approach utilizes open-source LLMs with fewer than 10 billion parameters, suitable for running on commodity hardware.
HySem achieves these requirements through a structured pipeline composed of three components: Context Optimizer Subsystem ($\mathcal{A}_{\text{CO}}$), Semantic Synthesizer ($\mathcal{A}_{\text{SS}}$), Syntax Validator ($\mathcal{A}_{\text{SV}}$). $\mathcal{A}_{\text{CO}}$ employs a novel pre-processing methodology for optimizing the context window utilized by HTML tables, enabling processing of large tables. $\mathcal{A}_{\text{SS}}$ utilises an LLM for transforming this optimized HTML table into semantic JSON. $\mathcal{A}_{\text{SV}}$ dissects the generated JSON for any syntax errors and outputs a syntactically-valid JSON.
Figure 2: Encoder subsystem of the Context Optimizer that rewrites the input tokens to an optimal set of tokens.
2.1 Context Optimizer Subsystem
The limited context size of large language models significantly constrains their ability to effectively process extensive tables. Although models with extended context lengths can alleviate this issue, they may sacrifice accuracy and increase processing times due to the quadratic complexity of self-attention mechanisms Tay et al. (2022) Vaswani et al. (2017). Recent advancements in Transformer architectures, such as Linformer Wang et al. (2020) and Flash Attention techniques Dao et al. (2022), have shown improvements in memory and time complexity. However, the necessity to reduce sequence length remains a fundamental challenge.
Figure 3: An example figure illustrating a table optimized by the Context Optimizer. (a) shows the HTML table in its original form without optimization and (b) shows the optimized HTML table. The cell content "MIC/MBC ($\mu$g/mL)" for example, represented in its original form in Figure 3(a), tokenizes into: [’MIC’, ’/M’, ’BC’, ’ (’, ’$\mu$’, ’ g’, ’/mL’, ’)’] containing 8 tokens. After optimization, the same cell content transforms into 2 tokens, represented in Figure 3(b) as: [’MIC’, ’/M’], resulting in a substantial reduction in token count.
Significant token inefficiency can occur when there is a misalignment between the tokenizer’s vocabulary and the domain-specific terminology in the input text. For example, the term "Amoxycillin," a common medication in pharmaceutical literature, is not part of the Llama 3 tokenizer’s vocabulary. As a result, it is represented by multiple tokens, thereby reducing token efficiency. This issue is further illustrated with the term "Potassium Clavulanate," which tokenizes as (’Pot’, ’assium’, ’ Cl’, ’v’, ’aul’, ’an’, ’ate’), resulting in seven tokens. Table 1. illustrates several key terms used in the pharmaceutical industry and their corresponding tokens obtained from Llama 3 tokenizer.
We introduce "Context Optimizer", a novel token alignment technique that performs dynamic token pruning to address these limitations and enhance context utilization. The Context Optimizer is implemented in 2 phases: Encoding phase and Decoding phase.
2.1.1 Encoding Phase
In the encoding phase, we begin with a preprocessing step where we remove tags and attributes from the HTML tables that do not contribute to semantic understanding, such as those meant for styling. This is followed by a minification step to further reduce the input size. By eliminating non-essential elements, we streamline the input data, making it more manageable for subsequent processing. A tokenizing step tokenizes the contents of each cell in the table. A cell may be empty in which case there are no tokens. A non-empty cell may transform to one or more tokens.
The goal of encoding is to arrive at representing the cell contents using a minimal set of tokens while ensuring a unique representation for each distinct cell content. The encoded inputs result in a valid HTML table with cell contents rewritten to require fewer tokens. A mapping table is maintained to track the original cell content and its optimized counterpart, facilitating later restoration. The encoded input is processed by the downstream modules in the HySem pipeline, including the Semantic Synthesizer and the Syntax Validator. Figure 3., illustrates a sample HTML table before and after optimization.
Collision Resolution: In the process of minimizing token sequences for efficient encoding of each cell content, it is possible for two distinct cell contents to map to some overlapping token sequences. To ensure each cell content maintains a unique representation, our method includes a collision resolution mechanism. The cell contents "4/8" and "4/4" in Figure 3(a) tokenize into ["4", "/", "8"] and ["4", "/", "4"], respectively, with the token sequence ["4", "/"] overlapping in both. After collision resolution, "4/8" is represented by the token sequence, ["4", "/"] while "4/4" transforms into ["4", "/", "4"], thus ensuring a unique representation as depicted in Figure 3(b).
2.1.2 Decoding Phase
In the decoding step, the output from the downstream processing is decoded to restore the original lexicon used in the table. This decoding step reverses the earlier encoding which is present in the model generated JSON, reconstructing the output to match the original input while retaining the benefits of the token optimization. The resulting JSON is both semantically accurate and more efficiently processed.
A key feature of our Context Optimizer is its dynamic nature, wherein the mapping of input words to an optimized token sequence and the corresponding decoding process are dependent solely on the current input, without relying on any static mapping.
2.2 Semantic Synthesizer
HTML tables are highly varied and arbitrary in nature, featuring elements such as row spans, column-spans, multi-level headers, nested tables and diverse data types. These intricacies make it challenging to accurately extract hierarchical and relational structures. Motivated by the LLM’s capability to comprehend deep semantic relationships, HySem adopts the open-sourced Meta-Llama-3-8B-Base model and fine-tunes it with a manually labeled dataset for transforming HTML tables into semantic JSON.
Concretely, the Semantic Synthesizer $\mathcal{A}_{\text{SS}}$ accepts a HTML table $\mathcal{H}_{\text{i}}$ pre-processed by the Context Optimizer, as input and produces JSON $\mathcal{J}_{\text{i}}$ as the output in the same encoded semantic space as the pre-processed input HTML table. Our initial experiments indicate that adopting Prompt Engineering to achieve this transformation results in the generation of less accurate JSON representations, as evidenced by our intrinsic and extrinsic evaluations. We found that the LLM is highly sensitive to prompts and struggles to effectively capture the wide variety present in these tables. We identified several common failure patterns present in the JSONs generated by the LLM, which are cataloged as Failure Modes in Table 2.
Failure Mode Example
Lexical errors "Characteristics" in place of "Characteristic"
Missing words "123" instead of "123 (n=25)"
Missing entire rows An entire row under a subheading missed
Misaligned keys/values A cell value placed under a different parent key
[…]
Table 2: Common failure modes and their examples.
To address these failure modes, we fine-tune the LLM to generate JSON representations that accurately reflect the input HTML table.
2.2.1 Dataset
For our dataset, we require inputs as HTML tables and labels as semantic JSON. We utilize the following two open-sourced datasets for tabular HTML sources:
PubTabNet: PubTabNet Smock et al. (2022) is a large-scale dataset for image-based table recognition, containing over 568,000 images of tables from scientific papers along with their corresponding HTML annotations.
FinTabNet: FinTabNet Zheng et al. (2021) is a dataset specifically designed for recognizing tables in financial documents, containing over 112,000 tables. Each table instance annotation includes fields such as the HTML structure and bounding box coordinates, similar to PubTabNet.
For our purposes, we extract the HTML from these sources and manually curate annotations for these tables. We filter tables that fit within the LLM’s context length (8k context window for Llama3). We randomly select and pre-process 1,364 HTML tables from both PubTabNet and FinTabNet using our Context Optimizer and manually curate corresponding JSON annotations. We split the data into 756 training samples and 608 testing samples. This dataset is used to train the Semantic Synthesizer. We chose these datasets as they are public and they are reasonable representations of the Pharmaceutical and Finance verticals respectively.
We also created hand labeled datasets to address the custom needs of the industry using the proprietary customer supplied data. These proprietary datasets are used to fine-tune models that use Semantic Synthesizer LLM as the base model.
2.2.2 Fine-tuning
We fine-tune the LLM in two stages. In the first stage, we utilize the open-sourced datasets discussed in the previous section, for fine-tuning our model. In the second stage, we manually curate JSON annotations for 130 randomly chosen tables from our proprietary APQR dataset and further fine-tune our model.
2.3 Syntax Validator
Syntax errors in the LLM-generated JSON output render the table unusable for further table processing, such as analytics. Therefore, correcting syntax errors present in the JSON output is a critical functionality, particularly for facilitating automated workflows in industry settings. To address these challenges, we developed a Syntax Validator, based on reflective agentic framework.
Specifically, the Syntax Validator $\mathcal{A}_{\text{sv}}$ accepts a prompt $\mathcal{P}$ and a syntactically invalid JSON $\mathcal{J}_{\text{i}}$ as inputs and produces a syntactically valid JSON $\mathcal{J}_{\text{v}}$ through iterative refinement.
Through self-reflection, $\mathcal{A}_{\text{sv}}$ iteratively refines the JSON output until a syntactically valid result is achieved or the maximum number of iterations is reached. The algorithm for this process is detailed in Algorithm 1.
Algorithm 1 Syntax Validator
1:Syntactically invalid JSON $\mathcal{J}_{\text{invalid}}$ from LLM
2:Syntactically valid JSON $\mathcal{J}_{\text{valid}}$
3:$\mathcal{J}_{\text{current}}$leftarrow $\mathcal{J}_{\text{invalid}}$
4:iterations $\leftarrow$ 0
5:while syntax errors in $\mathcal{J}_{\text{current}}$ and iterations $<$ max_iterations do
6:     if $\mathcal{J}_{\text{current}}$ is syntactically invalid then
7:         $\mathcal{J}_{\text{current}}$leftarrow $\mathcal{A}_{\text{sv}}$($\mathcal{J}_{\text{current}}$)
8:     end if
9:     iterations $\leftarrow$ iterations + 1
10:end while
11:$\mathcal{J}_{\text{valid}}$leftarrow $\mathcal{J}_{\text{current}}$
12:return $\mathcal{J}_{\text{valid}}$
3 Evaluation Methodology
For the accurate transformation of HTML tables into semantic JSON, two key objectives must be satisfied. First, content preservation: all strings from the HTML table cells must exactly occur in the JSON output, ensuring no information is lost during the conversion. Second, semantic accuracy: the generated JSON must accurately represent the hierarchical and relational structure of the original HTML table.
Building on this foundation, we develop intrinsic and extrinsic evaluation methods to measure the content and semantic accuracy of the JSON produced by HySem.
3.1 Intrinsic Evaluation
In intrinsic evaluation, we assess the representation of content from HTML cells in the generated JSON. We parse the HTML using Beautiful Soup to extract set of all the cell contents present in the HTML input, denoted as $\mathcal{H}_{\text{set}}=\{c_{1},c_{2},\ldots,c_{m}\}$. Similarly, we take the set of all elements in the JSON, denoted as $\mathcal{J}_{\text{set}}=\{e_{1},e_{2},\ldots,e_{n}\}$
To evaluate if each content item $c_{i}$ from $\mathcal{H}_{\text{set}}$ is present in $\mathcal{J}_{\text{set}}$, we define an indicator function $I(c_{i}\in\mathcal{J}_{\text{set}})$:
$I(c_{i}\in\mathcal{J}_{\text{set}})=\left\{\begin{array}[]{ll}1&\text{if }c_{i}\in\mathcal{J}_{\text{set}},\\ 0&\text{otherwise.}\end{array}\right.$
The Intrinsic Score (ISC) is then computed as:
$\text{ISC}=\frac{\sum_{i=1}^{m}I(c_{i}\in\mathcal{J}_{\text{set}})}{m}.$
Here, $m$ denotes the cardinality of $\mathcal{H}_{\text{set}}$.
3.2 Extrinsic Evaluation
Extrinsic evaluation assesses the semantic structure of the JSON by evaluating its ability to answer targeted questions. This method avoids direct comparison to the ground truth JSON ($\mathcal{G}_{\text{t}}$), which can vary in representation. Instead, we systematically validate the structure by formulating questions that probe specific semantic elements: In order to form a minimalistic set of questions that validate the structure, we leverage the paths from the root node to each leaf node in the JSON. The number of paths corresponds to the number of leaf nodes. Let $\mathcal{P}$ denote the set of these paths. Formally, we define $\mathcal{P}$ as:
$\mathcal{P}=\{p_{i}\mid p_{i}=(n_{1},n_{2},\ldots,n_{k})\}$
Here, $(n_{1},n_{2},\ldots,n_{k})$ represents a sequence of nodes starting from the root node $n_{1}$ to a leaf node $n_{k}$ in the JSON structure. For each path $p_{i}\in\mathcal{P}$, we prompt an LLM ($\mathcal{M}_{\text{q}}$) with $\mathcal{G}_{\text{t}}$ and $p_{i}$ as inputs, instructing it to generate a single question $\mathcal{Q}_{\text{i}}$. This question is targeted so that the value at the leaf node of the path is the expected answer $\mathcal{K}_{\text{e}}$.
The question-answer pairs are evaluated using an Evaluator LLM ($\mathcal{M}_{\text{eval}}$). This model accepts HySem generated JSON $\mathcal{J}_{\text{p}}$, question $\mathcal{Q}_{\text{i}}$ and an expected answer $\mathcal{K}_{\text{e}}$ as inputs. $\mathcal{M}_{\text{eval}}$ predicts an answer $\mathcal{K}_{\text{p}}$ and compares its prediction with the expected answer $\mathcal{K}_{\text{e}}$ in a single pass.
$Pred(p_{i})=\left\{\begin{array}[]{ll}1&\text{if }\mathcal{K}_{\text{p}}=\mathcal{K}_{\text{e}},\\ 0&\text{otherwise.}\end{array}\right.$
The Extrinsic Score (ESC) is computed as:
$ESC=\frac{1}{|\mathcal{P}|}\sum_{p_{i}\in\mathcal{P}}\text{Pred}(p_{i}).$
4 Results and Ablation Study
We present the results of HySem following the first fine-tuning stage, utilizing our 608 testing samples manually curated from the open-sourced FinTabNet and PubTabNet datasets. This comparison highlights HySem’s performance after being fine-tuned on publicly available data, ensuring a fair and transparent benchmark against other models.
Model Intrinsic Score % Extrinsic Score %
Meta-Llama-3-8B-Instruct 66 84.42
Phi-3-Medium-128K-Instruct 56 82.39
GPT-4o 93.4 90.15
HySem 91.12 88.39
Table 3: Intrinsic and Extrinsic Evaluation results
We benchmarked our model against popular open-source models, Meta-Llama-3-8B-Instruct and Microsoft’s Phi-3-Medium-128K-Instruct, as well as GPT-4o. When compared against Llama 3, HySem acheives over 25% higher accuracy on the intrinsic score and 3.97% higher on the extrinsic score. On the same test dataset, Microsoft Phi-3 Medium model performed poorly on the intrinsic metrics and fared relatively better at 82.39% on the extrinsic benchmark.
Number of test samples 608
Number of tokens before optimization (A) 570330
Number of tokens after optimization (B) 233287
Average number of tokens before optimization 938
Average number of tokens after optimization 383
Token Efficiency(%) 59.09
Table 4: Token efficiency metrics
We evaluated the token reduction efficiency of the Context Optimizer. Our studies as indicated by Table 4., show impressive reductions in the token count resulting in a reduced LLM inference throughput. We define Token Efficiency as
Token $\text{Token Efficiency}=\left(1-\frac{B}{A}\right)\times 100$ (1)
The results show that HySem significantly outperforms both the open-source models that we evaluated and remains competitive with the industry-leading OpenAI GPT-4o in both intrinsic and extrinsic measures with a token efficiency of over 59%. Although our model trails GPT-4o by 2% in accuracy metrics, it offers substantial advantages such as on-premise deployment on commodity hardware, cost-effectiveness, and enhanced data privacy—non-negotiable requirements for many enterprises.
5 Related Works
Recent research has explored various uses of HTML to enhance NLP tasks. Aghajanyan et al. (2021) developed the HyperText Language Model (HTLM), which leverages HTML’s structural elements to improve tasks like zero-shot summarization and question answering. Their approach utilizes HTML for structured prompting and template-based guidance but is largely confined to standard NLP tasks and assumes predefined HTML structures, such as <title> elements.
Gur et al. (2023) focused on HTML understanding through tasks like semantic classification, description generation, and autonomous web navigation. They identified the context window length as a significant bottleneck, noting that even models with support for 1000+ tokens struggle with performance degradation when processing larger snippet sizes. This challenge underscores the difficulty of managing extensive HTML content within fixed context constraints.
Sui et al. (2024) tackled context length limitations by serializing tables into text formats. Their strategies include varying input formats and identifying critical values through self-augmentation. While these methods help fit table data within the model’s context window, serialization can lead to the loss of structural nuances and detailed information.
In contrast, our work is broadly categorized as a Table Transformation task, where we focus on converting arbitrary HTML tables into semantic JSON. By employing a novel Context Optimizer, we preserve the table’s structural integrity while enabling the processing of large tables. Unlike methods relying on predefined templates or serialization, our approach does not make assumptions about the HTML structure.
6 Conclusion
We successfully implemented the HySem LLM Pipeline, which generates semantic JSON output from HTML tables. The Context Optimizer significantly improved context utilization, enabling us to process larger tables. Currently, we are piloting our product with a well-known pharmaceutical enterprise that operates multiple production plants and serves a global customer base.
This model powers several downstream applications, including analytics from regulatory compliance documents and the automatic creation of these documents.
Our future work includes supporting even larger tables that span multiple pages, developing techniques to improve processing speed, and building custom models for other verticals.